OpenStudy (anonymous):

evaluate: ∫1/x dx; n=8 on (1,2)

OpenStudy (across):

\[\int\frac{1}{x}dx.\]What's confusing about this integral?

OpenStudy (anonymous):

if you integrate it, you get x^0... so it's lnx but then i'm lost on finding the area

OpenStudy (across):

That's correct.\[\int\frac{1}{x}dx=\ln|x|.\]Now, what does that n stand for?

OpenStudy (anonymous):

the number of intervals?

OpenStudy (across):

I suppose (1, 2) is the range in which these intervals are partitioned?

OpenStudy (anonymous):

yes

OpenStudy (across):

If that's the case, then to solve this problem, you have to make use of the limit definition of an integral.

OpenStudy (anonymous):

okay... how would i write that out?

OpenStudy (across):

\[\int_{a}^{b}f(x)=\lim_{n\to\infty}\sum_{i=1}^{n}f(c_{i})\Delta x_{i}.\]

OpenStudy (anonymous):

is there a way to do it without the greek E? we are on chapter 7.3 and haven't learned that... but we do know limits

OpenStudy (jamesj):

So for your problem you have eight partitions?

OpenStudy (jamesj):

Assuming so, then let x_i = 1 + i/8, i =1, 2, 3, ..., 8. Then you want an approximation of the Riemann integral such as 1/8 [ f(x_1) + f(x_2) + ... + f(x_8) ] where f(x) = 1/x